When 127.68 g of brass at 163 degrees C are added to 150.00g of H2o at 22.4 degrees C, the final temperature of the brass and water is 32.6 degrees C. Calculate the value of the specific heat, c, for the brass.
1
Expert's answer
2018-02-07T15:16:08-0500
For water: q = cmΔT = 4.18 ∙ 150 (32.6 – 22.4) = 6395.4 J
For brass: c = q / mΔT = 6395 / 127.68 (163 – 32.6) = 0.384 J / gC